to take a technology course, Manu borrows $20 000 at an interest rate of 4.75% per annum compounded annunally. He plans to pay back the loan in 2 years. How much will Manu owe?


1
Expert's answer
2021-12-16T17:36:23-0500

"P=\$20000, r=4.75\%=0.0475, n=2 , A=?"


"A=P(1+r)^n=20000(1+0.0475)^2=\$21945.125"